Fermi National Laboratory


Accelerator Update for March 25, 2005 - March 28, 2005

Friday March 25
The day shift began with Operations monitoring store 4058 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam. NuMI is off due to a cooling problem with its target.

At 10:51 AM, a Booster RF station (BRF16) tripped off. It reset okay.

At 11:34 AM, the MiniBooNE horn tripped off for the first of two times in five minutes. It reset okay both times, but operators contacted an expert anyway.

At 1 PM, a BRF16 tripped off. It reset okay.

At 1:01 PM, Operations reported a problem with proton batches falling out of the Main Injector (MI) and then falling out on their way to Pbar. Operators notified an expert.

Operations terminated store 4058 at 1:38 PM.

Operations disabled beam to NuMI and Switchyard (SY) and put the TeV into shot setup at 1:55 PM.

Operations established store 4059 at 3:51 PM with an initial luminosity of 92.82E30. Pbar resumed taking beam.

Operations established beam to MCenter at 4:15 PM.

At 5:49 PM, BRF16 tripped off. It reset okay.

Operations established beam to Meson MTest experiment at 6:18 PM.

NuMI experts worked out a plan to resume taking beam in order to research their target problem. At 6:46 PM, Operations began sending low intensity beam to NuMI.

At 8:46 PM, BRF16 tripped off for the first of three times in eighty minutes. It reset okay each time.

Operations put the Recycler into shot setup to stash antiprotons at 10:45 PM.

The Recycler took 6 transfers of beam. Pbar resumed stacking at 11:50 PM.

Saturday March 26
The midnight shift began with Operations monitoring store 4059 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MCenter, and MTest taking beam.

At 2:31 AM, a Linac RF station (KRF2) tripped off on a LCW flow fault. An operator tapped on the flow switch and it reset okay.

At 2:54 AM, BRF2 tripped off twice. The second time it tripped operators investigated and had beam back up five minutes later.

At 3:30 AM, BRF16 tripped off for the first of three times in two hours. It reset okay each time.

At 10:16 AM, BRF16 tripped off. It reset okay.

At 1:24 PM, BRF16 tripped off for the first of two times in about thirty minutes. It reset okay both times.

At 3:07 PM, Operations reported that Pbar stopped stacking for about twelve minutes due to an emittance blow up. Later it was discovered that an expert was conducting a study.

At 4:16 PM, Operations reported that the 400 MeV area tripped off for the first of four times in about three hours.

At 4:20 PM, BRF16 tripped off. It reset okay.

Operations terminated store 4059 at 5:02 PM and put the TeV into shot setup.

At 6:32 PM, Operations reported a problem with the TeV chromaticity program. They rebooted a sector F0 scope, but it didn't help. Operations contacted an expert.

NuMI experts reported at 1:35 PM that they were removing shielding to get access to their target.

Operations established store 4060 at 8:07 PM, with an initial luminosity of 72.84E30. Pbar resumed stacking.

At 8:39 PM, Operations reported that the 400 MeV area tripped off for the first of five times in four hours. An expert came in to investigate.

At 10:23 PM, BRF16 tripped off. It reset okay.

At 11:53 PM, the MI permit tripped off due to MECAR. It reset okay.

Sunday March 27
The midnight shift began with Operations monitoring store 4060, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

At 12:09 AM, KRF2 tripped off on water flow. After some operator persuasion, it reset.

Operations put the Recycler into shot setup at 1:18 AM, but the transfers had problems. First, there were some trips during a reverse proton tune up. Then there was a trim magnet power supply (HV100) that wouldn't ramp and another trim (HT906) that wanted a large tune change that messed up efficiencies. During all this, an operator reported seeing a possible LCW leak below KRF6, which held off beam a short time.

At 2:57 AM, Operations reported that during the Recycler stashes they lost one of the transfers.

At 3:17 AM, BRF16 tripped off. It reset okay.

MCenter resumed taking beam at 3:50 AM.

After resolving a problem with M:HV100, Operations resumed sending beam to Pbar.

At 5:04 AM, BRF16 tripped off for the first of two times in forty-five minutes. It reset okay both times.

At 7 AM, a Linac expert turned off KRF6 to repair the LCW leak. Thirty-three minutes later he had it back up and operating. Pbar resumed stacking and MCenter resumed taking beam.

At 8:21 AM, the MiniBooNE horn tripped off. It reset okay.

At 9:28 AM, the 400 MeV area tripped off. It reset okay.

At 9:44 AM, BRF16 tripped off. It reset okay.

At 12:44 PM, Operation reported that the stack rate was dropping. Operators made changes that increase the rate for a while, but then it went into decline again. At 3:12 PM, an operator noticed that the target wasn't rotating. This resolved the problem.

At 3:38 PM, the Booster began having vacuum problems that held off beam. An expert came in and finished his repairs by 5:50 PM. Pbar resumed stacking.

Operations established beam to MCenter at 5:51 PM.

Operations terminated store 4060 at 8:03 PM and had the TeV in shot setup by 8:26 PM.

At 9:22 PM, a TeV quadrupole magnet power supply tripped off due to a quench indication. There was no quench and it reset okay.

Monday March 28
The midnight shift began with TeV in shot setup, with Operations monitoring the stash, and with MiniBooNE taking beam.

Operations established store 4062 at 12:25 AM with an initial luminosity of 94.7E30. Pbar resumed stacking.

MCenter resumed taking beam at 12:47 AM.

At 3:16 AM, the 400 MeV area tripped off. It reset okay.

At 4:34 AM, BRF13 tripped off and would not reset. Operators bypassed the station.

At 6:14 AM, BRF15 tripped off. It reset okay.

At 6:54 AM, KRF5 tripped off due to a water flow problem. It reset okay.

Operations put the Recycler into shot setup to stash antiprotons at 7:46 AM.

The Plan for Monday
The plan for today is to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/28/2005   email Fermilab